Jenny wants to utilize a portion of her garden for growing flowers. Being a designer herself, she visualizes a triangular area with different lengths for each side. For this Jenny decides to keep a side of the triangular area 5 feet shorter than the third side and the other side 4 feet longer than the third side. She also wants to restrict the perimeter of the triangular area to a maximum of 29 feet, so that it does not cover a huge area of the garden. Determine the length of each side of the triangular area that Jenny has visualized.
Answer by Earlsdon(6103)

Let the three sides of the triangle be A, B, and C, then:
A = C-5 "One side of the triangle is 5 ft. shorter than the third side.
B = C+4 "The other side is 4 ft. longer than the third side.", and...
A+B+C = 29 ft. "...restrict the perimeter...to a maximum of 29 ft...."
Substitute the first two equations for A and B into the third equation, to get:
(C-5)+(C+4)+C = 29 Simplify.
3C-1 = 29 Add 1 to both sides.
3C = 30 Divide both sides by 3.
C = 10ft.
A = C-5
A = 5ft.
B = C+4
B = 14ft.
The sides of the triangle would be:
5ft, 14ft,and 10ft.